A paper airplane follows a parabolic path with h=-1/4t^2+t+3, where h is height in meters, and t is time in seconds. Algebraically determine how long it takes for the paper airplane to hit the ground.
Answer by nerdybill(7384)

h=-1/4t^2+t+3
.
When the ball hits the ground, the height is zero.
So, set h to 0 and solve for t:
.
h=-1/4t^2+t+3
0 = -1/4t^2+t+3
multiply both sides by -4:
0 = t^2-4t-12
factoring:
0 = (t-6)(t+2)
t = {-2, 6}
.
Tossing out the negative solution:
t = 6 seconds